# Causality in the maximally extended Reissner–Nordström spacetime with identifications

### v1

@@ Abstract
The maximally extended Reissner–Nordström (RN) spacetime with $e^{2}<m^{2}$ can be interpreted either as an infinite chain of asymptotically flat regions connected by tunnels between timelike singularities or as a set of just one asymptotically flat region and one tunnel; the repetitions of this set in the infinite chain being identified. The second interpretation gives rise to the suspicion of acausality, i.e. the possibility of sending messages to one’s own past. A numerical investigation of this problem was carried out in this paper and gave the following result. Let E be the initial point of a radial timelike future-directed ingoing geodesic G, lying halfway between the outer horizon and the image of the null infinity in the maximally extended RN spacetime. Let E be the first future copy of E. It was verified whether the turning point of G will lie to the future or to the past from the past light cone (PLC) of E. In the second case the breach of causality does occur. It turned out that the acausality is present when $V_{E}$, the timelike coordinate of E, is negative with a sufficiently large $|V_{E}|$, and is absent with a sufficiently large $V_{E}>0$. In between these values there exists a $\widetilde{V}_{E}$, dependent on the initial data for the geodesic, for which the turning point lies on the PLC. So, the identification does lead to acausality. Nonradial timelike and null geodesics were also investigated, and a few hitherto unknown properties of the maximal extension were revealed. For example, the singularity arc at $r=0$ may be convex or concave, depending on the values of $m$ and $e$.
@@ 1. Motivation and summary
The maximally extended Reissner [1] – Nordström [2] (RN) spacetime with $e^{2}<m^{2}$ can be interpreted either as an infinite chain of asymptotically flat regions connected by tunnels between timelike singularities or as a set of just one asymptotically flat region and one tunnel; the repetitions of this set in the infinite chain being identified. The identification may be suspected of leading to acausality (i.e. an observer could supposedly send a message to its own past by means of timelike or null geodesics). A radial null geodesic sent into the tunnel will hit the singularity and will not get out into the next asymptotically flat region unless it is reflected somewhere in the tunnel. The problem was to verify what happens with timelike geodesics, radial and nonradial, and with nonradial null geodesics.
@@ 2. Basic facts about the maximally extended Reissner–Nordström spacetime
From (2.21) ${\rm e}^{2\gamma r^{*}}(r_{i})=0$ when $\gamma=\gamma_{i}$. Thus, from (2.14), the set $r=r_{i}$, in the coordinates that make it nonsingular, has the equation $u=\pm v$. Again by (2.21), ${\rm e}^{2\gamma r^{*}}(0)=$ constant $>0$. Hence, in the part of the $(u,v)$ coordinate plane wheree (2.14) applies, the equation of the singular set $r=0$ is $u^{2}-v^{2}=$ constant $>0$, i.e., it is a pair of hyperbolae that intersect the $u$-axis and are convex towards smaller $|u|$. However, where (2.16) applies, the constant $u^{2}-v^{2}$ is negative, and the image of the singular set is a pair of hyperbolae that intersect the $v$ axis. See Sec. 5. for more on this.
@@
We can proceed from a point E in the $r>r_{+}$ region back in time along a $q=$ constant null geodesic and cross the spurious singularity $r=r_{+}$ at $p=0$, or to the future along a $p=$ constant null geodesic and cross $r=r_{+}$ at $q=0$. By extending these two kinds of null geodesics, we cover sectors I, II and IV of Fig. 3.444This figure first appeared in the papers by Carter, see in particular Ref. [5]. By sending null geodesics back in time from sector II and to the future from sector IV we cover sector III. In the $r_{-}<r<r_{+}$ regions, we transform the $(u,v)$ coordinates continued from sector I to such $(u^{\prime},v^{\prime})$ that cancel the $r=r_{-}$ spurious singularity. The following remark will require attention in the next sections:
@@
Re ‘represent’: The arcs in Fig. 3 are an illustrative scheme, not an exact drawing. The shapes and positions in Fig. 3 of the exact images of lines of constant $r$ depend on the coordinates used and on the range of the $m$ and $e$ parameters. The schematic presentation thus avoids complications, but see Sec. 5. and Figs. 4, 5 and 6.
@@ 6. A timelike future-directed geodesic that does not break causality
G1a crosses $r=r_{+}$ smoothly using (3.11) – see the left panel of Fig. 8. As (2.9) shows, $\lim_{r\to r_{-}}r^{*}=+\infty$, $\lim_{r\to r_{+}}r^{*}=-\infty$ and ${{\rm d}{r^{*}}}/{{\rm d}{r}}=1/\phi<0$ for $r_{-}<r<r_{+}$, so with $\gamma=\gamma_{1}$ the function $r^{*}(r)$ is monotonic in this range and $r$ at $P_{2}$ is uniquely determined.
Fig. 8: Left panel: The segment G1a of geodesic $G_{1}$ crossing the $r=r_{+}$ horizon. Right panel: The segment G1b of $G_{1}$ passes through point $P_{4}$ – the image of $P_{2}$ under the transformation (4.6) – (4.7).
@@
4. We integrate (4.8) off $P_{3}$ with $\sigma=-1$. Between $r(P_{1})$ and $r(P_{2})$, G1a and G1b run side by side. The right panel of Fig. 8 shows that $P_{4}$ (the image of $P_{2}$ under the transformation (4.6) – (4.7)) lies on G1b with good precision.
@@ 7. The past light cone of E
As can be seen in Fig. 10, point $P_{10}$ lies well on the G2b segment.
Fig. 10: Closeup view on the neighbourhood of point $P_{10}$ in Fig. 4: it lies on G2b with good precision.
@@ 8. A nonradial timelike geodesic
Also the image of the second segment of $G_{J}$ (from the neighbourhood of $P_{3}$ to the neighbourhood of $P_{5}$) coincides with that of G1b at the scale of Fig. 4, and the images of their TPs are indistinguishable. The images of these segments are seen as separate only after magnification – see Fig. 12. The left panel shows the image of the neighbourhood of their initial points, the right panel shows the images of their final (turning) points. The ’old $P_{3}$’ is the image of the ’old $P_{1}$’ from Fig. 11 under the transformation (4.6) – (4.7), the ’new $P_{3}$’ is the image of the ’new $P_{1}$’ under the same transformation. Incidentally, Figs. 11 and 12 show that the transformation does not preserve the ordering of events along the ${\cal U}$ coordinate axis: ’new $P_{1}$’ is at a larger ${\cal U}$ than ’old $P_{1}$’, but ’new $P_{3}$’ is at a smaller ${\cal U}$ than ’old $P_{3}$. The two panels of Fig. 12 show that somewhere between $P_{3}$ and $P_{5}$ the image of $G_{J}$ intersects G1b, so $G_{J}$ goes at a (slightly) larger inclination to the ${\cal U}$ axis than G1b.
@@ 9. A nonradial null geodesic
Equations (8.3) and (8.4) still apply with $E_{J}$ replaced by $E_{N}$, and it is still true that with $\Gamma^{2}/E_{N}<1$ we have $r_{\rm 1tp}>r_{+}$. With $\Gamma^{2}/E_{N}>1$ it is still true that $r_{\rm 2tp}<r_{-}$. However, it is not always true that $r_{\rm 2tp}>r_{\rm tp}$ for the $r_{\rm tp}$ of (3.10), because here $(1-1/E_{N})$ may have any sign depending on the value of $r$, so (F2) does not imply (F3) in Appendix F.
@@ 11. Conclusions and summary
The main purpose of this paper was to verify whether an observer in an asymptotically flat region of the maximally extended Reissner – Nordström (RN) spacetime with $e^{2}<m^{2}$ can send a message (by means of timelike or null geodesics) to the past of his/her first future copy (see Figs. 3, 4 and 13). If this is possible, then the identifications of the asymptotically flat regions do lead to acausality. The answer is that geodesics emitted sufficiently late do not break causality, while those emitted sufficiently early do. Let the initial point of the geodesic be at event E, and let the first future copy of E be E. Then any radial null geodesic emitted at E will hit the singularity and stop there. For late-emitted timelike geodesics the turning point (TP) in the tunnel between the singularities lies to the future of the past light cone (PLC) of E, so the signal may come back to the observer only later than it was emitted. But for early-emitted timelike geodesics, the opposite happens: the TP lies to the past of the PLC of E. In this case, the geodesic continued beyond the TP can cross the worldline of E in the past of E, so the breach of causality is possible. The border between ’early’ and ’late’ are those emission points for which the TP lies right on the PLC of E.
@@
$\left(\Gamma^{2}-1\right)\ \left(-m+\sqrt{m^{2}-e^{2}+e^{2}\Gamma^{2}/E_{J}(r_{\rm 2tp})}\right)$ (F2)
$\leq\left(\Gamma^{2}/E_{J}(r_{\rm 2tp})-1\right)\ \left(-m+\sqrt{m^{2}-e^{2}+e^{2}\Gamma^{2}}\right).$
Both sides above are positive, so they may be squared and the direction of the inequality will not change. After squaring, using (F2) to eliminate

### v2

@@ Abstract
The maximally extended Reissner–Nordström (RN) spacetime with $e^{2}<m^{2}$ can be interpreted either as an infinite chain of asymptotically flat regions connected by tunnels between timelike singularities or as a set of just one pair of asymptotically flat regions and one tunnel; the repetitions of this set in the infinite chain being identified. The second interpretation gives rise to the suspicion of acausality, i.e. the possibility of sending messages to one’s own past. A numerical investigation of this problem was carried out in this paper and gave the following result. Let E be the initial point of a radial timelike future-directed ingoing geodesic G, lying halfway between the outer horizon and the image of the null infinity in the maximally extended RN spacetime. Let E be the first future copy of E. It was verified whether the turning point of G lies outside or inside the past light cone (PLC) of E. In the second case the breach of causality does occur. It turned out that the acausality is present when $V_{E}$, the timelike coordinate of E, is negative with a sufficiently large $|V_{E}|$, and is absent with a sufficiently large $V_{E}>0$. In between these values there exists a $\widetilde{V}_{E}$, dependent on the initial data for the geodesic, for which the turning point lies on the PLC. So, the identification does lead to acausality. Nonradial timelike and null geodesics were also investigated, and a few hitherto unknown properties of the maximal extension were revealed. For example, the singularity arc at $r=0$ may be convex or concave, depending on the values of $m$ and $e$.
@@ 1. Motivation and summary
The maximally extended Reissner [1] – Nordström [2] (RN) spacetime with $e^{2}<m^{2}$ can be interpreted either as an infinite chain of asymptotically flat regions connected by tunnels between timelike singularities or as a set of just one pair of asymptotically flat regions and one tunnel; the repetitions of this set in the infinite chain being identified. The identification may be suspected of leading to acausality (i.e. an observer could supposedly send a message to its own past by means of timelike or null geodesics). A radial null geodesic sent into the tunnel will hit the singularity and will not get out into the next asymptotically flat region unless it is reflected somewhere in the tunnel. The problem was to verify what happens with timelike geodesics, radial and nonradial, and with nonradial null geodesics.
@@ 2. Basic facts about the maximally extended Reissner–Nordström spacetime
From (2.21) ${\rm e}^{2\gamma r^{*}}(r_{i})=0$ when $\gamma=\gamma_{i}$. Thus, from (2.14), the set $r=r_{i}$, in the coordinates that make it nonsingular, has the equation $u=\pm v$. Again by (2.21), ${\rm e}^{2\gamma r^{*}}(0)=$ constant $>0$. Hence, in the part of the $(u,v)$ coordinate plane where (2.14) applies, the equation of the singular set $r=0$ is $u^{2}-v^{2}=$ constant $>0$, i.e., it is a pair of hyperbolae that intersect the $u$-axis and are convex towards smaller $|u|$. However, where (2.16) applies, the constant $u^{2}-v^{2}$ is negative, and the image of the singular set is a pair of hyperbolae that intersect the $v$ axis. See Sec. 5. for more on this.
@@
We can proceed from a point E in the $r>r_{+}$ region back in time along a $q=$ constant null geodesic and cross the spurious singularity $r=r_{+}$ at $p=0$, or to the future along a $p=$ constant null geodesic and cross $r=r_{+}$ at $q=0$. By extending these two kinds of null geodesics, we cover sectors I, II and IV of Fig. 3.444This figure first appeared in the papers by Carter, see in particular Ref. [5]. By sending null geodesics back in time from sector II and to the future from sector IV we cover sector III. In the $r_{-}<r<r_{+}$ regions, we transform the $(u,v)$ coordinates continued from sector I to such $(u^{\prime},v^{\prime})$ that cancel the $r=r_{-}$ spurious singularity. The following remark will require attention in the subsequent sections:
@@
Re the meaning of ‘represent’: The arcs in Fig. 3 are an illustrative scheme, not an exact image. The shapes and positions in Fig. 3 of the exact images of lines of constant $r$ depend on the coordinates used and on the range of the $m$ and $e$ parameters. The schematic presentation thus avoids complications, but see Sec. 5. and Figs. 4, 5 and 6.
@@ 6. A timelike future-directed geodesic that does not break causality
G1a crosses $r=r_{+}$ smoothly using (3.11). As (2.9) shows, $\lim_{r\to r_{-}}r^{*}=+\infty$, $\lim_{r\to r_{+}}r^{*}$ =absent=$\infty$= - ∞ and ${{\rm d}{r^{*}}}/{{\rm d}{r}}=1/\phi<0$ for $r_{-}<r<r_{+}$, so with $\gamma=\gamma_{1}$ the function $r^{*}(r)$ is monotonic in this range and $r$ at $P_{2}$ is uniquely determined.
@@
4. We integrate (4.8) off $P_{3}$ with $\sigma=-1$. Between $r(P_{1})$ and $r(P_{2})$, G1a and G1b run side by side.
@@ 8. A nonradial timelike geodesic
Also the image of the second segment of $G_{J}$ (from the neighbourhood of $P_{3}$ to the neighbourhood of $P_{5}$) coincides with that of G1b at the scale of Fig. 4, and the images of their TPs are indistinguishable. The images of these segments are seen as separate only after magnification – see Fig. 10. The left panel shows the image of the neighbourhood of their initial points, the right panel shows the images of their final (turning) points. The ’old $P_{3}$’ is the image of the ’old $P_{1}$’ from Fig. 9 under the transformation (4.6) – (4.7), the ’new $P_{3}$’ is the image of the ’new $P_{1}$’ under the same transformation. Incidentally, Figs. 9 and 10 show that the transformation does not preserve the ordering of events along the ${\cal U}$ coordinate axis: ’new $P_{1}$’ is at a larger ${\cal U}$ than ’old $P_{1}$’, but ’new $P_{3}$’ is at a smaller ${\cal U}$ than ’old $P_{3}$. The two panels of Fig. 10 show that somewhere between $P_{3}$ and $P_{5}$ the image of $G_{J}$ intersects G1b, so $G_{J}$ goes at a (slightly) larger inclination to the ${\cal U}$ axis than G1b.
@@ 9. A nonradial null geodesic
Equations (8.3) and (8.4) still apply with $E_{J}$ replaced by $E_{N}$, and it is still true that with $\Gamma^{2}/E_{N}<1$ we have $r_{\rm 1tp}>r_{+}$. With $\Gamma^{2}/E_{N}>1$ it is still true that $r_{\rm 2tp}<r_{-}$. However, it is not always true that $r_{\rm 2tp}>r_{\rm tp}$ for the $r_{\rm tp}$ of (3.10), because here $(1-1/E_{N})$ may have any sign depending on the value of $r$, so (F) does not imply (F3) in Appendix F.
@@ 11. Conclusions and summary
The main purpose of this paper was to verify whether an observer in an asymptotically flat region of the maximally extended Reissner – Nordström (RN) spacetime with $e^{2}<m^{2}$ can send a message (by means of timelike or null geodesics) to the past of his/her first future copy (see Figs. 3, 4 and 11). If this is possible, then the identifications of the asymptotically flat regions do lead to acausality. The answer is that geodesics emitted sufficiently late do not break causality, while those emitted sufficiently early do. Let the initial point of the geodesic be at event E, and let the first future copy of E be E. Then any radial null geodesic emitted at E will hit the singularity and stop there. For late-emitted timelike geodesics the turning point (TP) in the tunnel between the singularities lies to the future of the past light cone (PLC) of E, so the signal may come back to the observer only later than it was emitted. But for early-emitted timelike geodesics, the opposite happens: the TP lies to the past of the PLC of E. In this case, the geodesic continued beyond the TP can cross the worldline of E in the past of E, so the breach of causality is possible. The border between ’early’ and ’late’ are those emission points for which the TP lies right on the PLC of E.
@@
$\left(\Gamma^{2}-1\right)\ \left(-m+\sqrt{m^{2}-e^{2}+e^{2}\Gamma^{2}/E_{J}(r_{\rm 2tp})}\right)$
$\leq\left(\Gamma^{2}/E_{J}(r_{\rm 2tp})-1\right)\ \left(-m+\sqrt{m^{2}-e^{2}+e^{2}\Gamma^{2}}\right).$ (F2)
Both sides above are positive, so they may be squared and the direction of the inequality will not change. After squaring, using (F) to eliminate